what is the present value of Rs 10,000 due in 2 years at 8%p.a


1
Expert's answer
2021-11-10T15:32:41-0500

"A=10,000"

"n=2"

"i=0.08"

Using the yearly means

"A=P(1+i)^n"

"10,000=P(1+0.08)^2"

"10,000=P(1.08)^2"

"10,000=1.1664P"

"P=8,573.38"